Drew is making a cardboard cut out of a right triangle. He knows that the longest side must be 15 inches long. One of the short sides must be 8 inches. How many inches long will he make the other side? Use the appropriate formula to find the solution and round to the nearest tenth. Show all steps or reasoning used.
So far, I have no idea what I'm supposed to do.
I've determined that since it's a right triangle the angle is 90 degrees. The long side is the diagonal side that would connect the vertical side and the base together which is 15 inches and the base is 8inches long. I don't know where to go from here though.

Answer by stanbon(75887)

Drew is making a cardboard cut out of a right triangle. He knows that the longest side must be 15 inches long. One of the short sides must be 8 inches. How many inches long will he make the other side? Use the appropriate formula to find the solution and round to the nearest tenth. Show all steps or reasoning used.
----
hypotenuse = 15 inches
One of the legs = 8 inches
Let the other leg = x
============================
Equation:
Use Pythagoras to solve for "x":
x^2 + 8^2 = 15^2
x^2 = 225-64
x = sqrt(161)
x = 12.69 inches
-----